OpenStudy (anonymous):

A circle is defined by the following equation: x^(2)+y^(2)-8x+14y+47=0 What are the coordinates of its center?

OpenStudy (anonymous):

(4,-7) :P

OpenStudy (anonymous):

How did you get that so fast?

OpenStudy (anonymous):

magic :p

OpenStudy (anonymous):

Okay magician, a circle is defined by the following equation: 3x^(2)+3y^(2)+30x+30y+150=0 What is its radius?

OpenStudy (anonymous):

now the radius takes a little longer lol

OpenStudy (anonymous):

x^2 +Y^2 +10X + 10Y + 50 =0 X^2 +10X+25 +Y^2 + 10Y + 25 + 50 =0 (X+5)^2 +(Y+5)^2 +50 =0

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

should be (x+5)^2 +(y+5)^2 =0, WHICH IS NOT A CIRCLE

OpenStudy (anonymous):

So then is 0 not the answer?

OpenStudy (anonymous):

how can u have a circle of zero radius :P

OpenStudy (anonymous):

a small one :)

OpenStudy (anonymous):

idk, i have this as a qustion. it doesn't make sence.

OpenStudy (anonymous):

teacher made a mistake

OpenStudy (anonymous):

It's a computer program though.

OpenStudy (anonymous):

well, its not a circle , a sum of sqaures cannot be zero ( in the real numbers anyway )

OpenStudy (anonymous):

well, its not a circle , a sum of sqaures cannot be zero ( in the real numbers anyway )

OpenStudy (anonymous):

the answer is definately not zero

OpenStudy (anonymous):

would you say it is 10?

OpenStudy (anonymous):

no I would say its not a circle